# DyRoNet: Dynamic Routing and Low-Rank Adapters for Autonomous Driving Streaming Perception

### v3

@@ 1 Introduction
In autonomous driving systems, it is crucial to achieve low-latency and high-precision perception. Traditional object detection algorithms Zou et al. (2023), while effective in various contexts, often confront the challenge of latency due to inherent computational delays. This lag between algorithmic processing and real-world states can lead to notable discrepancies between predicted and actual object locations. Such latency issues have been extensively reported and are known to significantly impact the decision-making process in autonomous driving systems Chen et al. (2023).
Addressing these challenges, the concept of streaming perception has been introduced as a response Li et al. (2020). This perception task aims to predict “future” results by accounting for the delays incurred during the frame processing stage. Unlike traditional methods that primarily focus on detection at a given moment, streaming perception transcends this limitation by anticipating future environmental states, and aligning perceptual outputs closer to real-time dynamics. This new paradigm is key in addressing the critical gap between real-time processing and real-world changes, thereby enhancing the safety and reliability of autonomous driving systems Muhammad et al. (2020).
Figure 1: Illustration of DyRoNet’s dynamic selection mechanism in streaming perception. This diagram showcases DyRoNet’s capability to adaptively choose the most suitable perception strategy, contrasting with the static approach of traditional methods in complex environments [Viewing in color and at an expanded scale].
Although the existing streaming approach seems promising, it still faces contradictions in real-world scenarios. These contradictions primarily stem from the diverse and unpredictable nature of driving environments. The factors such as camera motion, weather conditions, lighting variations, and the presence of small objects seriously impact the performance of perception measures, leading to fluctuations that challenge their robustness and reliability (see Sec. 3.1). This complexity in real-world scenarios underscores the limitations of a single, uniform model, which often struggles to adapt to the varied demands of different driving conditions Guo et al. (2019). In general, the challenges of streaming perception mainly include:
@@
Generally speaking, these challenges highlight the demand for streaming perception. As we study in Sec. 3.1, addressing the diverse scenario distribution and achieving an optimal balance between performance and efficiency are key to advancing the state-of-the-art in autonomous driving. To address the intricate challenges presented by real-world streaming perception, we introduce DyRoNet, a framework designed to enhance dynamic routing capabilities in autonomous driving systems. DyRoNet stands as a low-rank enhanced dynamic routing framework, specifically crafted to cater to the requirements of streaming perception. It encapsulates a suite of pre-trained branch networks, each meticulously fine-tuned to optimally function under distinct environmental conditions. A key component of DyRoNet is the speed router module, ingeniously developed to assess and efficiently route input data to the most appropriate branch, as detailed in Sec. 3.2. To sum up, the contributions are listed as:
@@
  • By utilizing a variety of sophisticated streaming perception techniques, DyRoNet provides the speed router as a major invention. This component dynamically determines the best route for handling each input, ensuring efficiency and accuracy in perception. The ability to adapt and be versatile is demonstrated by this dynamic route-choosing mechanism.
@@
In summary, DyRoNet offers advancements for low-latency, high-accuracy perception in autonomous driving. By addressing challenges of environmental adaptability and dynamic branch selection, DyRoNet sets new benchmarks in achieving low-latency and high-accuracy perception.
Figure 2: The DyRoNet Framework: This figure presents the architecture of DyRoNet, featuring a multi-branch network design. For simplicity, only two branches are shown, each representing a streaming perception sub-network. The core network architecture is detailed in the upper right. Each branch processes both the current frame $I_{t}$ and a series of historical frames $I_{t-1},I_{t-2},\cdots,I_{t-n}$. The backbone and neck of the network extract features, which are then split into two streams for the current and historical frames. These streams are fused together before entering the prediction head. Branch selection is governed by the Speed Router, which analyzes the frame difference $\Delta I_{t}$ derived from $I_{t}$ and $I_{t-1}$ to determine the most suitable branch for the given input.
@@ 2.1 Streaming Perception
The existing streaming perception methods fall into three main categories. (1) The initial methods focused on single-frame, with models like YOLOv5 Jocher et al. (2021) and YOLOX Ge et al. (2021) achieving real-time performance. However, lacking motion trend capture, they struggle in dynamic scenarios. (2) The recent approaches incorporated current and historical frames, like StreamYOLO Yang et al. (2022) building on YOLOX with dual-flow fusion. LongShortNet Li et al. (2023) used longer histories and diverse fusion. DAMO-StreamNet He et al. (2023) added asymmetric distillation and deformable convolutions to improve large object perception. (3) Recognizing the limitations of single models, current methods explore dynamic multi-model systems. One approach Ghosh et al. (2021) adapts models to environments via reinforcement learning. DaDe Jo et al. (2022) extends StreamYOLO by calculating delays to determine frame steps. A later version Huang and Chen (2023) added multi-branch prediction heads. Beyond 2D detection, streaming perception expands into optical flow, tracking, and 3D detection, with innovations in metrics and benchmarks Wang et al. (2023c); Sela et al. (2022); Wang et al. (2023b). Distinct from these existing approaches, our proposed method, DyRoNet, introduces a low-rank enhanced dynamic routing mechanism specifically designed for streaming perception. DyRoNet stands out by integrating a suite of advanced branch networks, each fine-tuned for specific environmental conditions. Its key innovation lies in the speed router module, which not only routes input data efficiently but also dynamically adapts to the diverse and unpredictable nature of real-world driving scenarios.
@@ 2.2 Dynamic Neural Networks
Dynamic Neural Networks (DNNs) feature adaptive network selection, outperforming static models in efficiency and performance (Han et al. (2021); Lan et al. (2023); Zhang et al. (2023)). The existing research primarily focuses on structural design for core deep learning tasks like image classification (Huang et al. (2018); Wang et al. (2020, 2018)). DNNs follow two approaches: (1) Multi-branch models (Bejnordi et al. (2019); Cai et al. (2021); Shazeer et al. (2017); Wang et al. (2023a); Qiao et al. (2022)) rely on a lightweight router assessing inputs to direct them to appropriate branches, enabling tailored computation. (2) By generating new weights based on inputs (Yang et al. (2019); Chen et al. (2020); Su et al. (2019); Zhu et al. (2019)), these models dynamically alter computations to match diverse needs. DNN applications expand beyond conventional tasks. In object detection, DynamicDet (Lin et al. (2023)) categorizes inputs and processes them through distinct branches. This illustrates DNNs’ broader applicability and efficiency, promising contributions particularly for complex, dynamic environments.
@@ 3.1 Motivation for DyRoNet
Analysis of Influential Factors. Our statistical analysis of the Argoverse-HD dataset Li et al. (2020) underscores the profound influence of environmental dynamics on the effectiveness of streaming perception. While weather inconsistently impacts accuracy, suggesting the presence of other influential factors (see Appendix A.1), fluctuations in the object count show limited correlation with performance degradation (see Appendix A.2). Conversely, the presence of small objects across various scenes poses a significant challenge for detection, especially under varying motion states (see Appendix A.3). Notably, disparities in performance are most pronounced across different environmental motion states (see Appendix A.4), thereby motivating the need for a dynamic, velocity-aware routing mechanism in DyRoNet.
Rationale for Dynamic Routing. Analysis reveals that StreamYOLO’s reliance on a single historical frame falters at high velocities, in contrast to multi-frame models, highlighting a clear connection between velocity and detection performance (see Tab. 1). Dynamic adaptation of frame history, based on vehicular speed changes, enables DyRoNet to strike a balance between accuracy and latency (see Sec. 4.3). Through first-order differences, the system efficiently switches models to align with environmental motions. Specifically, the dynamic routing is designed to select the optimal architecture based on the vehicle’s speed profile, ensuring precision at lower velocities for detailed perception and efficiency at higher speeds for swift response. Such adaptable routing, informed by comprehensive speed analysis, positions DyRoNet as a robust solution for reliable perception across diverse autonomous driving scenarios. Next, we introduce DyRoNet in detail.
@@ 3.2 Architecture of DyRoNet
Mathematically, the core of DyRoNet lies the processing of a frame sequence, $\mathcal{S}=\{I_{t},\cdots,I_{t-N\delta t}\}$, where $N$ indicates the number of frames and $\delta t$ the interval between successive frames. The process of the framework is formalized as:
@@
Correspondingly, the weights of DyRoNet are denoted by $\mathcal{W}=\{W^{d},W^{l},W^{r}\}$, where $W^{d}$ indicates the weights of the streaming perception model, $W^{l}$ relates to the Low-Rank Adaptation (LoRA) weights within each model, and $W^{r}$ pertains to the Router Network. The culmination of this process is the final output, $\mathcal{T}$, a compilation of feature maps. These maps can be further decoded through $Decode(\mathcal{T})$, revealing essential details like objects, categories, and locations. Below we introduce each module in detail.
Figure 3: The mean curves of frame differences are depicted here. The four curves correspond to frame sizes of the original frame, 200$\times$200, 100$\times$100, and 50$\times$50. Notably, these curves show distinct fluctuations across different vehicle motion scenarios.
Router Network. The Router Network in DyRoNet plays a crucial role in understanding and classifying the dynamics of the environment. This module is designed for both environmental classification and branch decision-making. To effectively and rapidly capture environmental speed, frame differences are employed as the input to the Router Network. As shown in Fig. 3, frame differences exhibit a high discriminative advantage for different environmental speeds.
@@
$\mathcal{T}$ $=\mathcal{F}(\mathcal{S},\mathcal{P},W)$
$=\text{Disp}(\mathcal{R}(\Delta I_{t}),\mathcal{P})(I_{t};W^{d}_{\sigma},W^{l}_{\sigma})$
where $\mathcal{R}$ symbolizes the Router Network, and $\Delta I_{t}$ refers to the frame difference, a key input for model selection. The weights $W^{d}_{\sigma}$ and $W^{l}_{\sigma}$ correspond to the selected streaming perception model and its Low-Rank Adaptation (LoRA) parameters, respectively.
Note that the versatility of DyRoNet is further highlighted by its compatibility with a wide range of Streaming Perception models, even ones that rely solely on detectors Ge et al. (2021). To demonstrate the efficacy of DyRoNet, it has been evaluated using three contemporary streaming perception models: StreamYOLO Yang et al. (2022), LongShortNet Li et al. (2023), and DAMO-StreamNet He et al. (2023) (see Sec. 4.3). This Model Bank & Dispatcher strategy illustrates the adaptability and robustness of DyRoNet across different streaming perception scenarios.
@@
To address these challenges, we have incorporated the Low-Rank Adapter Hu et al. (2021) into our streaming perception models. Within each model $P_{i}$, initially pre-trained on a dataset, the key components are the convolution kernel and bias matrices, symbolized as $W^{d}_{i}$. The rank of the Low-Rank Adaptation (LoRA) module is defined as $r$, a value significantly smaller than the dimensionality of $W^{d}_{i}$, to ensure efficient adaptation. The update to the weight matrix adheres to a low-rank decomposition form, represented as $W_{d}^{i}+\delta W=W_{d}^{i}+BA$ [^2] Here, $B$ is a matrix in $R^{d\times r}$, and $A$ is in $R^{r\times k}$, ensuring that the rank $r$ remains much smaller than $d$. This adaptation strategy allows for the original weights $W_{d}^{i}$ to remain fixed, while the low-rank components $BA$ are trained and adjusted. The adaptation process is executed through the following projection:
@@
where $x$ represents the input image or feature map, and $\Delta W=W^{l}_{i}=BA$. The matrices $A$ and $B$ start from an initialized state and are fine-tuned during the adaptation process. This approach maintains the general applicability of the model by fixing $W_{d}^{i}$, while also enabling specialization within specific sub-domains, as determined by Router Network.
Particularly, in DyRoNet, we employ a rank $r$ of $32$ for the LoRA module, though this can be adjusted based on specific requirements of the scenarios in question. This low-rank adaptation mechanism not only enhances the flexibility of the DyRoNet framework but also significantly mitigates the risk of overfitting, ensuring that each branch remains efficient and effective in its designated role.
@@ 3.3 Training Details of DyRoNet
The training process of DyRoNet focuses on two primary goals: (1) improving the performance of individual branches within the streaming perception model and (2) achieving an optimal balance between accuracy and computational efficiency. This dual-objective framework is represented by the overall loss function:
@@
[…]
@@
[…]
@@
[…]
@@
[…]

### v4

@@ 1 Introduction
In autonomous driving systems, it is crucial to achieve low-latency and high-precision perception. Traditional object detection algorithms [40], while effective in various contexts, often confront the challenge of latency due to inherent computational delays. This lag between algorithmic processing and real-world states can lead to notable discrepancies between predicted and actual object locations. Such latency issues have been extensively reported and are known to significantly impact the decision-making process in autonomous driving systems [4].
Addressing these challenges, the concept of streaming perception has been introduced as a response [20]. This perception task aims to predict “future” results by accounting for the delays incurred during the frame processing stage. Unlike traditional methods that primarily focus on detection at a given moment, streaming perception transcends this limitation by anticipating future environmental states, and aligning perceptual outputs closer to real-time dynamics. This new paradigm is key in addressing the critical gap between real-time processing and real-world changes, thereby enhancing the safety and reliability of autonomous driving systems [23].
Figure 1: Illustration of DyRoNet’s adaptive selection mechanism in streaming perception, contrasting with static traditional methods in complex environments [Best viewed in color and enlarged].
Although the existing streaming approach seems promising, it still faces contradictions in real-world scenarios. These contradictions primarily stem from the diverse and unpredictable nature of driving environments. The factors such as camera motion, weather conditions, lighting variations, and the presence of small objects seriously impact the performance of perception measures, leading to fluctuations that challenge their robustness and reliability (see Sec. 3.1). This complexity in real-world scenarios underscores the limitations of a single, uniform model, which often struggles to adapt to the varied demands of different driving conditions [8]. In general, the challenges of streaming perception mainly include:
@@
Generally speaking, these challenges highlight the demand for streaming perception. As we study in Sec. 3.1, addressing the diverse scenario distribution and achieving an optimal balance between performance and efficiency are key to advancing the state-of-the-art in autonomous driving. To address the intricate challenges presented by real-world streaming perception, we introduce DyRoNet, a framework designed to enhance dynamic routing capabilities in autonomous driving systems. DyRoNet stands as a low-rank enhanced dynamic routing framework, specifically crafted to cater to the requirements of streaming perception. It encapsulates a suite of pre-trained branch networks, each meticulously fine-tuned to optimally function under distinct environmental conditions. A key component of DyRoNet is the speed router module, ingeniously developed to assess and efficiently route input to the optimal branch, as detailed in Sec. 3.2. To sum up, the contributions are listed as:
@@
  • By utilizing a variety of streaming perception techniques, DyRoNet provides the speed router as a major invention. This component dynamically determines the best route for handling each input, ensuring efficiency and accuracy in perception. The ability to adapt and be versatile is demonstrated by this dynamic route-choosing mechanism.
@@
Figure 2: The DyRoNet Framework: This figure shows DyRoNet’s architecture with a multi-branch network. Two branches are illustrated, each as a streaming perception sub-network. The upper right details the core architecture. Each branch processes the current frame $I_{t}$ and historical frames $I_{t-1},I_{t-2},\cdots,I_{t-n}$. Features are extracted by the backbone and neck, split into streams for current and historical frames, fused, then passed to the prediction head. The Speed Router selects the branch based on frame difference $\Delta I_{t}$ from $I_{t}$ and $I_{t-1}$.
@@ 2.1 Streaming Perception
The existing streaming perception methods fall into three main categories. (1) The initial methods focused on single-frame, with models like YOLOv5 [15] and YOLOX [6] achieving real-time performance. However, lacking motion trend capture, they struggle in dynamic scenarios. (2) The recent approaches incorporated current and historical frames, like StreamYOLO [35] building on YOLOX with dual-flow fusion. LongShortNet [19] used longer histories and diverse fusion. DAMO-StreamNet [10] added asymmetric distillation and deformable convolutions to improve large object perception. (3) Recognizing the limitations of single models, current methods explore dynamic multi-model systems. One approach [7] adapts models to environments via reinforcement learning. DaDe [14] extends StreamYOLO by calculating delays to determine frame steps. A later version [13] added multi-branch prediction heads. Beyond 2D detection, streaming perception expands into optical flow, tracking, and 3D detection, with innovations in metrics and benchmarks [32, 25, 30]. Distinct from these existing approaches, our proposed method, DyRoNet, introduces a low-rank enhanced dynamic routing mechanism specifically designed for streaming perception. DyRoNet stands out by integrating a suite of advanced branch networks, each fine-tuned for specific environmental conditions. Its key innovation lies in the speed router module, which not only routes input data efficiently but also dynamically adapts to the diverse and unpredictable nature of real-world driving scenarios.
@@ 2.2 Dynamic Neural Networks
Dynamic Neural Networks (DNNs) feature adaptive network selection, outperforming static models in efficiency and performance [9, 17, 36]. The existing research primarily focuses on structural design for core deep learning tasks like image classification [12, 31, 29]. DNNs follow two approaches: (1) Multi-branch models [1, 3, 26, 28, 24, 33, 18] rely on a lightweight router assessing inputs to direct them to appropriate branches, enabling tailored computation. (2) By generating new weights based on inputs [34, 5, 27, 39], these models dynamically alter computations to match diverse needs. DNN applications expand beyond conventional tasks. In object detection, DynamicDet [21] categorizes inputs and processes them through distinct branches. This illustrates DNNs’ broader applicability and potential for dynamic environments.
@@ 3.1 Motivation for DyRoNet
Analysis of Influential Factors. Statistical analysis of the Argoverse-HD dataset [20] underscores the profound influence of environmental dynamics on the effectiveness of streaming perception. While weather inconsistently impacts accuracy, suggesting the presence of other influential factors (see Appendix A.1), fluctuations in the object count show limited correlation with performance degradation (see Appendix A.2). Conversely, the presence of small objects across various scenes poses a significant challenge for detection, especially under varying motion states (see Appendix A.3). Notably, disparities in performance are most pronounced across different environmental motion states (see Appendix A.4), thereby motivating the need for a dynamic, velocity-aware routing mechanism in DyRoNet.
Rationale for Dynamic Routing. Analysis reveals that StreamYOLO’s reliance on a single historical frame falters at high velocities, in contrast to multi-frame models, highlighting a connection between speed and detection performance (see Tab. 2). Dynamic adaptation of frame history, based on vehicular speed, enables DyRoNet to strike a balance between accuracy and latency (see Sec. 4.3). Through first-order differences, the system efficiently switches models to align with environmental motions. Specifically, the dynamic routing is designed to select the optimal architecture based on the vehicle’s speed profile, ensuring precision at lower velocities for detailed perception and efficiency at higher speeds for swift response. These comprehensive analysis imforms DyRoNet as a robust solution for reliable perception across diverse autonomous driving scenarios.
@@ 3.2 Architecture of DyRoNet
Mathematically, the core of DyRoNet lies the processing of a frame sequence, $\mathcal{S}=\{I_{t},\cdots,I_{t-N\delta t}\}$, where $N$ indicates the number of frames and $\delta t$ the interval between successive frames. The framework process is formalized as:
@@
Correspondingly, the weights of DyRoNet are denoted by $\mathcal{W}=\{W^{d},W^{l},W^{r}\}$, where $W^{d}$ indicates the weights of the streaming perception model, $W^{l}$ relates to the Low-Rank Adaptation (LoRA) weights within each model, and $W^{r}$ pertains to the Router Network. The culmination of this process is the final output, $\mathcal{T}$, a compilation of feature maps. These maps can be further decoded through $Decode(\mathcal{T})$, revealing essential details like objects, categories, and locations.
Figure 3: The mean curves of frame differences are depicted here. The four curves correspond to frame sizes of the original frame, 200$\times$200, 100$\times$100, and 50$\times$50. Notably, these curves show distinct fluctuations across different vehicle motion cases.
Router Network. The Router Network in DyRoNet plays a crucial role in understanding and classifying the dynamics of the environment. This module is designed for both environmental classification and branch decision-making. To effectively capture environmental speed, frame differences are employed as the input to the Router Network. As shown in Fig. 3, frame differences exhibit a high discriminative advantage for different environmental speeds.
@@
$\mathcal{T}=\mathcal{F}(\mathcal{S},\mathcal{P},W)=\text{Disp}(\mathcal{R}(\Delta I_{t}),\mathcal{P})(I_{t};W^{d}_{\sigma},W^{l}_{\sigma}),$
where $\mathcal{R}$ symbolizes the Router Network, and $\Delta I_{t}$ refers to the frame difference, a key input for model selection. The weights $W^{d}_{\sigma}$ and $W^{l}_{\sigma}$ correspond to the selected streaming perception model and its LoRA parameters, respectively.
Note that the versatility of DyRoNet is further highlighted by its compatibility with a wide range of Streaming Perception models, even ones that rely solely on detectors [6]. To demonstrate the efficacy of DyRoNet, it has been evaluated using three contemporary streaming perception models: StreamYOLO [35], LongShortNet [19], and DAMO-StreamNet [10] (see Sec. 4.3). This Model Bank & Dispatcher strategy illustrates the adaptability and robustness of DyRoNet across different streaming perception scenarios.
@@
To address these challenges, we have incorporated the LoRA [11] into DyRoNet. Within each model $P_{i}$, initially pre-trained on a dataset, the key components are the convolution kernel and bias matrices, symbolized as $W^{d}_{i}$. The rank of the LoRA module is defined as $r$, a value significantly smaller than the dimensionality of $W^{d}_{i}$, to ensure efficient adaptation. The update to the weight matrix adheres to a low-rank decomposition form, represented as $W_{d}^{i}+\delta W=W_{d}^{i}+BA$ [^2] Here, $B$ is a matrix in $R^{d\times r}$, and $A$ is in $R^{r\times k}$, ensuring that the rank $r$ remains much smaller than $d$. This adaptation strategy allows for the original weights $W_{d}^{i}$ to remain fixed, while the low-rank components $BA$ are trained and adjusted. The adaptation process is executed through the following projection:
@@
where $x$ represents the input image or feature map, and $\Delta W=W^{l}_{i}=BA$. The matrices $A$ and $B$ start from an initialized state and are fine-tuned during adaptation. This approach maintains the general applicability of the model by fixing $W_{d}^{i}$, while also enabling specialization within specific sub-domains, as determined by Router Network.
In DyRoNet, we employ $r=32$ for the LoRA module, though this can be adjusted based on specific requirements of the scenarios in question. This low-rank adaptation mechanism not only enhances the flexibility of the DyRoNet framework but also mitigates the risk of overfitting, ensuring that each branch remains efficient and effective in its designated role.
@@ 3.3 Training Details of DyRoNet
The training process of DyRoNet focuses on two primary goals: (1) Improving the performance of individual branches. The backpropagation only updates the chosen model’s weights in this step, enabling fine-tuning on segregated samples. (2) Achieving an optimal balance between accuracy and computational efficiency. This step only train the speed router while the remaining branches are frozen. This dual-objective framework is represented by the overall loss function:
@@
where $\mathcal{L}^{sp}$ represents the streaming perception loss, and $\mathcal{L}^{E^{2}}$ denotes the effective and efficient (E2) loss, which supervises branch selection.
@@
Effective and Efficient (E2) Loss. During the training phase, streaming perception loss values from all branches are compiled into a vector $v^{sp}\in\mathbb{R}^{K}$, and inference time costs are aggregated into $v^{time}\in\mathbb{R}^{K}$, with $K$ indicating the total number of branches in DyRoNet. To account for hardware variability, a normalized inference time vector $\hat{v}^{time}=\mathrm{softmax}(v^{time})$ is introduced. This vector is derived using the Softmax function to minimize the influence of hardware discrepancies. The representation for effective and efficient (E2) decision-making is defined as:
@@
where $\mathcal{O}$ denotes one-hot encoding, producing a boolean vector of length $K$, with the value of $1$ at the index representing the estimated optimal branch at that moment. The E2 Loss is then formulated as:
@@
DyRoNet and relevant techniques. Although the structure of DyRoNet is somewhat similar to MoE[26, 37], in contrast, the gate network of MoE is not well-suited for streaming perception. This limitation has led to the development of speed router and the $\mathcal{L}^{E^{2}}$ loss function. DyRoNet, for efficiency, chooses one model over MoE’s multiple experts, aligning with MoE in concept but differing in gate structure and selection strategy, making it unique for streaming contexts.
The speed router is inspired by Network Architecture Search (NAS). However, it uniquely addresses the challenge of streaming perception by transforming the search problem into an optimization of coded distances. The formulation of the loss function, $\mathcal{L}^{E^{2}}$, involves converting $v^{\text{time}}$ into a distribution via softmax, which is then combined with $v^{\text{sp}}$ to determine the optimal model through an one-hot vector $f^{E^{2}}$. This approach effectively simplifies the intricate problem of balancing accuracy and latency into a more tractable optimization task. Instead of employing NAS for loss search, our design is intricately linked to the specific needs of streaming perception, with KL divergence selected for its robustness in noisy situations[16]. This demonstrates the efficiency and innovation of our approach.
Overall, the process of training DyRoNet involves striking a meticulous balance between the SP loss, which ensures the efficacy of each branch, and the E2 loss, which optimizes efficiency. The primary objective of this training is to develop a model that not only delivers high accuracy in perception tasks but also operates within acceptable latency constraints, which is a critical requirement for real-time applications. This balanced approach enables DyRoNet to adapt dynamically to varying computational resources and environmental conditions, thereby maintaining optimal performance in diverse streaming perception scenarios.